OpenStudy (anonymous):

I don't want the answer, just need to know how to do it? Help? "What is the sum of a 52–term arithmetic sequence where the first term is 1 and the last term is 154?"

OpenStudy (valpey):

You could work out what the common difference is, then use the finite sequence sum formula. \[\large s_n = \frac{n(t_1 + t_n)}{2}\]

OpenStudy (valpey):

Actually, I guess you could skip the first step. Sn is the sum of n terms. t_1 is the first term and t_n is the nth term. n is the number of terms.

OpenStudy (anonymous):

So would n=154?

OpenStudy (anonymous):

OH! so x=154(155)/2?

OpenStudy (valpey):

No. It says there are 52 terms so n = 52.

OpenStudy (valpey):

So you aren't adding 1+2+3+4+...154. You are adding \[1+t_2+t_3+t_4+... + t_{49}+t_{50}+t_{51}+154\]

OpenStudy (anonymous):

So...if I'm correct. (52/2)(154+1)?

OpenStudy (anonymous):

then, 26*155? =4030?

OpenStudy (valpey):

good.